explain the pressure correction in Wander Wall correction.


(p+an^2/v^2).


shouldn't it be (p-an^2/v^2) 

Grade:

1 Answers

Akash Sood
12 years ago

No it is (P+an^2/V^2) because in case real gas molecules are attracted by other molecules inthe container, so the pressure appear on the wall's of container is less than expected pressure. So some additional pressure is required so as to make it a pressure equal to the pressure of an ideal gas.

P correct=P observed + p

P is directly proportional to d^2

d=n/V

therefore, d is directly proportional to n^2/V^2 i.e, p=an^2/V^2

a is coefficient of attraction.

 

Don't forget to approve my answer!!